Not in text book.
Can you work this equation so i can see the steps you took in working it?
I worked this equation and came up with Y-10 but it does not check. What step am i doing wrong?
7y-2(3-y)=4(y=1)
This is how i worked it
7y - 2(3-y)=4(y+1)
7y - 6 - 2y= 4y+4
5y - 6 =4y+4
4y 4y
1y - 6 =+4
+6 +6
y=10
What am i dong wrong? Thanks i really appreciate your help

Answer by Hook(36)

This is how i worked it
7y - 2(3-y)=4(y+1) <==========on the left side, -2*-y = +2y
7y - 6 - 2y= 4y+4
5y - 6 =4y+4
4y 4y
1y - 6 =+4
+6 +6
y=10